OpenStudy (anonymous):

how do i do this?x + y = 3 Which of the following equations is equivalent to the given equation? @gypsy1274

OpenStudy (anonymous):

sorry it was 1/2x + 1/3y = 3

OpenStudy (anonymous):

I don't understand the question. Am I missing something?

OpenStudy (anonymous):

is this equivalent to 1/2x+1/3y=3? 2x + 3y = 18

OpenStudy (anonymous):

No, they are not equivalent.

OpenStudy (anonymous):

18/6 = 3 but 2/6 = 1/3 not 1/2

OpenStudy (anonymous):

3x + 2y = 3 2x + 3y = 18 3x + 2y = 18 are any of these equivalent to it if so explain how you know please.

OpenStudy (anonymous):

you still there? @gypsy1274

OpenStudy (anonymous):

Yes, sorry. CHecking....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which is the original equation?

OpenStudy (anonymous):

1/2x + 1/3y = 3

OpenStudy (anonymous):

Multiply through by 6...

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

is it always multilpy by 6 for theese problems? @gypsy1274 ?

OpenStudy (anonymous):

No, it could be any number. But you do have to distribute the number through the entire equation. I chose 6 because 3*6=18

OpenStudy (anonymous):

ok thanks again

OpenStudy (anonymous):

Your Welcome.